# A note on a cluster structure of the coordinate ring of a simple algebraic group

### v2

@@ 1. Introduction
In the foundational paper [FZ02] of cluster algebras, Fomin and Zelevinsky conjectured that the coordinate rings of several varieties related to a simply-connected simple algebraic group $G$ have natural structures of cluster algebra. This conjecture has been widely investigated and verified. See, for example, [FZ02, FZ03, BFZ05, Sco06, GLS08, GLS11, Wil13, Lec16, SSBW19, FWZ, GY21, SW21, GY23, SSB24, CGGLSS25].
As for the coordinate ring $\mathbb{C}[G]$ of $G$ itself, $\mathbb{C}[SL_{2}]$ is mentioned as a typical example of cluster algebra in [FZ02]. In [FWZ], it is shown that $\mathbb{C}[SL_{n}]$ is a cluster algebra for $n\geq 2$. However, in the literature, there does not seem to be a proof of the fact that $\mathbb{C}[G]$ has a cluster structure for a general simply-connected simple algebraic group $G$. The goal of this paper is to provide a proof of this fact in the case when $G$ is not of type $F_{4}$. The precise statement is the following.
@@
In [BFZ05, Theorem 2.10], it is shown that $\mathscr{U}^{\circ}$ coincides with the coordinate ring $\mathbb{C}[G^{w_{0},w_{0}}]$ of the open double Bruhat cell $G^{w_{0},w_{0}}$ in $G$, and in [GY23], [SW21, Theorem 4.13], it is shown that $\mathscr{U}^{\circ}=\mathscr{A}^{\circ}$, including the case of type $F_{4}$. Here $\mathscr{A}^{\circ}$ and $\mathscr{U}^{\circ}$ denote the cluster and upper cluster algebras obtained from $\mathscr{A}$ and $\mathscr{U}$, respectively, by adding the inverses of frozen variables. However, the fact that $\mathbb{C}[G^{w_{0},w_{0}}]=\mathscr{A}^{\circ}=\mathscr{U}^{\circ}$ does not immediately imply $\mathbb{C}[G]=\mathscr{A}=\mathscr{U}$. Indeed, all the equalities $\mathbb{C}[G]=\mathscr{A}$, $\mathscr{A}=\mathscr{U}$, and $\mathbb{C}[G]=\mathscr{U}$ seem to be non-trivial in general, while the inclusion $\mathscr{A}\subset\mathscr{U}$ follows from the general fact, known as the Laurent phenomenon, of the theory of cluster algebras ([FZ02, Theorem 3.1], [FZ03, Proposition 11.2]). The delicate relation between $\mathscr{A}^{\circ}=\mathscr{U}^{\circ}$ and $\mathscr{A}=\mathscr{U}$ is discussed in [BMS19]. The third equality $\mathbb{C}[G]=\mathscr{U}$ was announced in [Qin, Claim 7.6] for an arbitrary $G$ without proof [^1] While the author was preparing this paper, he was informed that a proof of $\mathbb{C}[G]=\mathscr{U}$ is given in [QY] using a different method. The author is grateful to Fan Qin and Milen Yakimov for letting him know about their ongoing work., and the proof is now provided in a very recent preprint [QY, Theorem B.1].
@@
In our proof of the main theorem, the result [GLS13, Theorem 1.4] (see Theorem 2.1) is essential. It reduces the proof of the equalities in the main theorem to the proof of the inclusion $\mathbb{C}[G]\subset\mathscr{A}$. Here we do not use the assumption that $G$ is not of type $F_{4}$. Then, under the assumption, we prove $\mathbb{C}[G]\subset\mathscr{A}$ by showing that $\mathbb{C}[G]$ is generated by the generalized minors. This fact is shown in [IOS23, Proposition 2.2] when $G$ is of type $A_{n}$, $B_{n}$, $C_{n}$, $D_{n}$, $E_{6}$ or $E_{7}$. Therefore, the main contribution of this paper is to show this property in the cases of type $E_{8}$ and $G_{2}$. The remark on our assumption concerning $F_{4}$ can be found in Remark 3.10.
@@ Acknowledgments
The author would like to thank Tsukasa Ishibashi and Linhui Shen for numerous helpful discussions and comments. He is thankful to Bernard Leclerc since a question from him served as a strong motivation for writing this paper. The author is grateful to Fan Qin and Milen Yakimov for the kind explanation of their works. The author is also grateful to Luca Francone and Bernard Leclerc for letting him know their ongoing works. The author would like to thank anonymous referees for reading the manuscript carefully and for suggesting many improvements.
This work was supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists Grant Number 23K12950.
@@ 2.1. Simple algebraic groups
We set $S=\{1,\dots,n\}$, following the labeling in [Kac, Chapter 4]. Let $\mathfrak{h}:=\sum_{s\in S}\mathbb{C}h_{s}$ be a Cartan subalgebra of $\mathfrak{g}$. Define the weight lattice $P$ and the monoid of dominant integral weights $P_{+}$ by
@@
$r_{s}(\mu)=\lambda-\langle h_{s},\mu\rangle\alpha_{s}$
@@
Let $G$ be a simply-connected simple algebraic group over $\mathbb{C}$ whose Lie algebra $\operatorname{Lie}(G)$ is isomorphic to $\mathfrak{g}$, and fix a maximal torus $H$ of $G$. We consider an isomorphism $\operatorname{Lie}(G)\xrightarrow[]{\sim}\mathfrak{g}$ which maps the Lie algebra of $H$ to $\mathfrak{h}$. For $s\in S$, we write the root subgroup of $G$ corresponding to $\alpha_{s}$ (resp. $-\alpha_{s}$) as $U_{s}^{+}$ (resp. $U_{s}^{-}$). We take one-parameter subgroups $x_{s}\colon\mathbb{C}\xrightarrow{\sim}U_{s}^{+}$ and $y_{s}\colon\mathbb{C}\xrightarrow{\sim}U_{s}^{-}$ so that there exists a homomorphism $\varphi_{s}\colon SL_{2}(\mathbb{C})\to G$ of algebraic groups, satisfying
@@
$\bigoplus_{\lambda\in P_{+}}V(\lambda)^{\ast}\boxtimes V(\lambda)\xrightarrow{\sim}\mathbb{C}[G],\ f\otimes v\mapsto C^{V(\lambda)}(f,v),$
which is an algebraic version of the Peter–Weyl theorem. See, for example, [GW, Theorem 4.2.7].
@@ 2.2. Cluster algebras
We have the inclusion $\mathscr{A}(\bm{x},B)\subset\mathscr{U}(\bm{x},B)$, which is known as the Laurent phenomenon ([FZ02, Theorem 3.1], [FZ03, Proposition 11.2]). The elements $x_{i}$ ($i\in J_{\mathrm{fr}}$) are called frozen variables. We remark that the frozen variables are always non-invertible in the cluster algebra that we deal with in this paper.
@@
The following is one of the main result of [GLS13].
@@ 2.3. Berenstein–Fomin–Zelevinsky seeds
Let $(u,v)\in W\times W$. In [BFZ05, Theorem 2.10], it is shown that there is a structure of an upper cluster algebra with invertible frozen variables on the coordinate ring $\mathbb{C}[G^{u,v}]$ of the double Bruhat cell $G^{u,v}=B^{+}\overline{u}B^{+}\cap B^{-}\overline{v}B^{-}$. Here $B^{+}$ and $B^{-}$ denote the Borel subgroups containing $H$ corresponding to $\Phi^{+}$ and $-\Phi^{+}$, respectively.
@@
$\bm{x}_{\bm{s}}=\{\Delta(k;\bm{s})\}_{k\in J}.$
Note that the set of frozen variables is equal to $\{\Delta_{\varpi_{s},w_{0}\varpi_{s}},\Delta_{w_{0}\varpi_{s},\varpi_{s}}\mid s\in S\}$.
@@
$\Delta(k;\bm{s})=\Delta_{w\varpi_{s},w^{\prime}\varpi_{s}}$
@@
and the same is true for $w^{\prime}$. Moreover, if $w=e$ (resp. $w^{\prime}=e$), there is no condition on $(s_{i_{1}},\dots,s_{i_{N/2}})$ (resp. $(s_{j_{1}},\dots,s_{j_{N/2}})$). Set
@@
Note that neither $w_{0}\varpi_{s}$ nor $w\varpi_{s}$ with $w\varpi_{s}<\varpi_{s}$ are dominant.
@@ 3.1. Main result
We do not state that the statements of Theorems 3.1 and 3.2 do not hold in the case when $G$ is of type $F_{4}$. Indeed, $\mathbb{C}[G]=\mathscr{U}$ is proved for an arbitrary $G$ in a very recent preprint [QY, Theorem B.1] by a different method.
@@
It suffices to show that $C^{V}(f,v)$ can be expressed as a polynomial in generalized minors for weight vectors $f\in(V^{\ast})_{\mu}$ and $v\in V_{\nu}$. If $\mu,\nu\in\operatorname{wt}(V)\setminus\{0\}$, then $C^{V}(f,v)$ is a generalized minor up to a constant. Hence we deal with the case when $\mu$ or $\nu$ equals 00.
@@
for some $v^{\prime}_{\eta}\in V_{\eta}$ and $v^{\prime\prime}_{-\eta}\in V_{-\eta}$. Hence, if we write
$m^{\ast}(f)=\sum_{i}f^{\prime}_{i}\otimes f^{\prime\prime}_{i},$
then
@@
$=C^{V\otimes V}\left(\sum\nolimits_{i}f^{\prime}_{i}\otimes f^{\prime\prime}_{i},\sum\nolimits_{\eta\in P;\eta\neq 0}v^{\prime}_{\eta}\otimes v^{\prime\prime}_{-\eta}\right)$
$=\sum_{i}\left(\sum_{\eta\in P;\eta\neq 0}C^{V}(f^{\prime}_{i},v^{\prime}_{\eta})C^{V}(f^{\prime\prime}_{i},v^{\prime\prime}_{-\eta})\right).$
@@
When $\mathfrak{g}$ is of type $E_{8}$ or $G_{2}$, there exists a quasi-minuscule $\mathfrak{g}$-module $V$ and a surjective homomorphism $m\colon V\otimes V\to V$ of $\mathfrak{g}$-modules satisfying $m(V_{0}\otimes V_{0})=\{0\}$.
@@
Let $\mathfrak{g}$ be a simple Lie algebra of type $E_{8}$ or $G_{2}$. Consider the quasi-minuscule $\mathfrak{g}$-module $V$ stated in Theorem 3.9. Then the corresponding representation of $G$ is a faithful representation $\rho\colon G\to GL(V)$ which gives a closed immersion $G\to SL(V)$, since the center of $G$ is trivial in this case (see [Bor, Lemma 22.1], [GW, Remark 11.2.16], [Bri10, Corollary 1.13]). Therefore, the coordinate ring $\mathbb{C}[G]$ is generated by the matrix coefficients of $V$ (see, for example, [IOS23, Proposition 2.1]), and hence, by the generalized minors thanks to Theorem 3.8. ∎
@@
[…]
When $\mathfrak{g}$ is of type $F_{4}$, there does not exist a homomorphism $m$ of $\mathfrak{g}$-modules satisfying the assumption in Theorem 3.8. Indeed, a quasi-minuscule $\mathfrak{g}$-module is isomorphic to the 26-dimensional irreducible module $V(\varpi_{4})$ ($\dim V(\varpi_{4})_{0}=2$). We have $V(\varpi_{4})\otimes V(\varpi_{4})\simeq V(2\varpi_{4})\oplus V(\varpi_{1})\oplus V(\varpi_{3})\oplus V(\varpi_{4})$. Therefore, a surjective homomorphism of $\mathfrak{g}$-modules $m\colon V(\varpi_{4})\otimes V(\varpi_{4})\to V(\varpi_{4})$ is uniquely determined up to constant multiplication, and it is the only candidate of the homomorphism $m$ satisfying the assumption in Theorem 3.8. However, we have $m(V(\varpi_{4})_{0}\otimes V(\varpi_{4})_{0})\neq\{0\}$. The Lie algebra $\mathfrak{g}$ of type $F_{4}$ is realized as the Lie algebra of derivations of the $27$-dimensional exceptional Jordan algebra $\mathfrak{J}$ over $\mathbb{C}$. See, for example, [Yok, Section 2] for its explicit presentation. Hence there is an action of $\mathfrak{g}$ on $\mathfrak{J}$, and $\mathfrak{J}$ is isomorphic to $V(\varpi_{4})\oplus V(0)$ as a $\mathfrak{g}$-module. Here the trivial $\mathfrak{g}$-module $V(0)$ corresponds to the space spanned by the unit $\mathbf{1}$ of $\mathfrak{J}$. Then the multiplication $m_{\mathfrak{J}}$ of $\mathfrak{J}$ induces a surjective homomorphism of $\mathfrak{g}$-modules $m\colon V(\varpi_{4})\otimes V(\varpi_{4})\to V(\varpi_{4})$ as follows:
[…]
However, we can directly check that $m(V(\varpi_{4})_{0}\otimes V(\varpi_{4})_{0})\neq\{0\}$ since the weight space $V(\varpi_{4})_{0}$ corresponds to the space spanned by the trace-free diagonal matrices in $\mathfrak{J}$ under the presentation in [Yok, Section 2]. See [Yok] for further details.

### v3

@@ 1. Introduction
In the foundational paper [FZ02] of cluster algebras, Fomin and Zelevinsky conjectured that the coordinate rings of several varieties related to a simply-connected simple algebraic group $G$ have natural structures of cluster algebra. This conjecture has been widely investigated and verified. See, for example, [FZ02, FZ03, BFZ05, Sco06, GLS08, GLS11, Wil13, Lec16, SSBW19, FWZ21, GY21, SW21, GY23, SSB24, CGGLSS25].
As for the coordinate ring $\mathbb{C}[G]$ of $G$ itself, $\mathbb{C}[SL_{2}]$ is mentioned as a typical example of cluster algebra in [FZ02]. In [FWZ21], it is shown that $\mathbb{C}[SL_{n}]$ is a cluster algebra for $n\geq 2$. However, in the literature, there does not seem to be a proof of the fact that $\mathbb{C}[G]$ has a cluster structure for a general simply-connected simple algebraic group $G$.
The goal of this paper is to provide a proof of this fact in the case when $G$ is not of type $F_{4}$. The precise statement is the following.
@@
In [BFZ05, Theorem 2.10], it is shown that $\mathscr{U}^{\circ}$ coincides with the coordinate ring $\mathbb{C}[G^{w_{0},w_{0}}]$ of the open double Bruhat cell $G^{w_{0},w_{0}}$ in $G$, and in [GY23], [SW21, Theorem 4.13], it is shown that $\mathscr{U}^{\circ}=\mathscr{A}^{\circ}$, including the case of type $F_{4}$. Here $\mathscr{A}^{\circ}$ and $\mathscr{U}^{\circ}$ denote the cluster and upper cluster algebras obtained from $\mathscr{A}$ and $\mathscr{U}$, respectively, by adding the inverses of the frozen variables. However, the fact that $\mathbb{C}[G^{w_{0},w_{0}}]=\mathscr{A}^{\circ}=\mathscr{U}^{\circ}$ does not immediately imply $\mathbb{C}[G]=\mathscr{A}=\mathscr{U}$. Indeed, all the equalities $\mathbb{C}[G]=\mathscr{A}$, $\mathscr{A}=\mathscr{U}$, and $\mathbb{C}[G]=\mathscr{U}$ seem to be non-trivial in general, while the inclusion $\mathscr{A}\subset\mathscr{U}$ follows from the general fact, known as the Laurent phenomenon, of the theory of cluster algebras ([FZ02, Theorem 3.1], [FZ03, Proposition 11.2]). The delicate relation between $\mathscr{A}^{\circ}=\mathscr{U}^{\circ}$ and $\mathscr{A}=\mathscr{U}$ is discussed in [BMS19]. The third equality $\mathbb{C}[G]=\mathscr{U}$ was announced in [Qin25, Claim 7.6] for an arbitrary $G$ without proof, and the proof is now provided in a very recent preprint [QY25, Theorem B.1] [^1] While the author was preparing this paper, he was informed that a proof of $\mathbb{C}[G]=\mathscr{U}$ is given in [QY25] using a different method. The author is grateful to Fan Qin and Milen Yakimov for letting him know about their ongoing work..
Here we make a remark on an application of the main result of this paper. When we study the algebraic structure on the ring of regular functions on some variety $X$ (over $\mathbb{C}$), a morphism $X\to G$ often plays an important role, since it induces an algebra homomorphism $\mathbb{C}[G]\to\mathbb{C}[X]$ via pullback. In such a situation, the cluster structure on $\mathbb{C}[G]$ is useful for studying the cluster structure on $\mathbb{C}[X]$. For example [^2] In a recent work [FL25], Francone and Leclerc study the cluster structure on the coordinate ring of the scheme $B(G,c)$, called $(G,c)$-bands ($c$ is a Coxeter element), by using the projection maps $B(G,c)\to G$ and the result of this paper., when $X$ is the moduli space $\mathcal{A}_{G,\Sigma}^{\times}$ of generically decorated twisted $G$-local systems on a marked surface $\Sigma$ with boundaries, one can define a family of morphisms $\mathcal{A}_{G,\Sigma}^{\times}\to G$, called Wilson lines [IOS23]. These morphisms are essentially used in the study of the cluster structure on $\mathbb{C}[\mathcal{A}_{G,\Sigma}^{\times}]$. Indeed, in [IOS23], $\mathbb{C}[G]=\mathscr{A}$ was implicitly proved in the case when $G$ is not of type $E_{8},F_{4},G_{2}$, and it is precisely the assumption on $G$ in [IOS23, Theorem 1]. By the main result of this paper, we can show that the result parallel to [IOS23, Theorem 1] holds also in the case of type $E_{8}$ and $G_{2}$. The detailed discussion will appear in a separate publication. Note that we cannot restrict the Wilson lines to the morphisms $\mathcal{A}_{G,\Sigma}^{\times}\to G^{w_{0},w_{0}}$, hence the corresponding morphisms $\mathbb{C}[G]\to\mathbb{C}[\mathcal{A}_{G,\Sigma}^{\times}]$ cannot be extended to $\mathbb{C}[G^{w_{0},w_{0}}]\to\mathbb{C}[\mathcal{A}_{G,\Sigma}^{\times}]$ in general.
@@
In our proof of the main theorem, the result [GLS13, Theorem 1.4] (see Theorem 2.1) is essential. It reduces the proof of the equalities in the main theorem to the proof of the inclusion $\mathbb{C}[G]\subset\mathscr{A}$. Here we do not use the assumption that $G$ is not of type $F_{4}$. Then, under the assumption, we prove $\mathbb{C}[G]\subset\mathscr{A}$ by showing that $\mathbb{C}[G]$ is generated by the generalized minors. This fact is shown in [IOS23, Proposition 2.2] when $G$ is of type $A_{n}$, $B_{n}$, $C_{n}$, $D_{n}$, $E_{6}$ or $E_{7}$. Therefore, the main contribution of this paper is to show this property in the cases of type $E_{8}$ and $G_{2}$. Remark 3.10 explains why the techniques developed here to deal with the $E_{8}$ and $G_{2}$ cases do not work for $F_{4}$.
@@ Acknowledgments
The author would like to thank Tsukasa Ishibashi and Linhui Shen for numerous helpful discussions and comments. He is greatly indebted to Bernard Leclerc, whose question served as a strong motivation for writing this paper. He also wishes to express his thanks to Fan Qin and Milen Yakimov for kindly explaining their works. He is grateful to Luca Francone and Bernard Leclerc for letting him know their ongoing works. He would like to thank the anonymous referees for their careful reading of the manuscript and for suggesting many improvements.
This work was supported by JSPS KAKENHI Grant-in-Aid for Early-Career Scientists, Grant Number 23K12950.
@@ 2.1. Simple algebraic groups
We set $S=\{1,\dots,n\}$, following the labeling in [Kac90, Chapter 4]. Let $\mathfrak{h}:=\sum_{s\in S}\mathbb{C}h_{s}$ be a Cartan subalgebra of $\mathfrak{g}$. Define the weight lattice $P$ and the monoid of dominant integral weights $P_{+}$ by
@@
$r_{s}(\mu)=\mu-\langle h_{s},\mu\rangle\alpha_{s}$
@@
Let $G$ be a simply-connected simple algebraic group over $\mathbb{C}$ whose Lie algebra $\operatorname{Lie}(G)$ is isomorphic to $\mathfrak{g}$, and fix a maximal torus $H$ of $G$. We consider an isomorphism $\operatorname{Lie}(G)\xrightarrow[]{\sim}\mathfrak{g}$ which maps the Lie algebra of $H$ to $\mathfrak{h}$. For $s\in S$, we write the root subgroup of $G$ corresponding to $\alpha_{s}$ (resp. $\alpha_{s}$) as $U_{s}^{+}$ (resp. $U_{s}^{-}$). We take additive one-parameter subgroups $x_{s}\colon\mathbb{C}\xrightarrow{\sim}U_{s}^{+}$ and $y_{s}\colon\mathbb{C}\xrightarrow{\sim}U_{s}^{-}$ so that there exists a homomorphism $\varphi_{s}\colon SL_{2}(\mathbb{C})\to G$ of algebraic groups, satisfying
@@
$\bigoplus_{\lambda\in P_{+}}V(\lambda)^{\ast}\boxtimes V(\lambda)\xrightarrow{\sim}\mathbb{C}[G],\ f\otimes v\mapsto C^{V(\lambda)}(f,v)\ (f\in V(\lambda)^{\ast},v\in V(\lambda)),$
which is an algebraic version of the Peter–Weyl theorem. See, for example, [GW09, Theorem 4.2.7].
@@ 2.2. Cluster algebras
We have the inclusion $\mathscr{A}(\bm{x},B)\subset\mathscr{U}(\bm{x},B)$, which is known as the Laurent phenomenon ([FZ02, Theorem 3.1], [FZ03, Proposition 11.2]). The elements $x_{i}$ ($i\in J_{\mathrm{fr}}$) are called frozen variables. We remark that the inverses of the frozen variables are not added in the definition of the cluster algebras in this paper.
@@
The following is one of the main results of [GLS13].
@@ 2.3. Berenstein–Fomin–Zelevinsky seeds
Let $(u,v)\in W\times W$. In [BFZ05, Theorem 2.10], it is shown that there is a structure of upper cluster algebra with invertible frozen variables on the coordinate ring $\mathbb{C}[G^{u,v}]$ of the double Bruhat cell $G^{u,v}=B^{+}\overline{u}B^{+}\cap B^{-}\overline{v}B^{-}$. Here $B^{+}$ and $B^{-}$ denote the Borel subgroups containing $H$ corresponding to $\Phi^{+}$ and $\Phi^{+}$, respectively.
@@
$\bm{x}_{\bm{s}}=(\Delta(k;\bm{s}))_{k\in J}.$
Note that the set of the frozen variables is $\{\Delta_{\varpi_{s},w_{0}\varpi_{s}},\Delta_{w_{0}\varpi_{s},\varpi_{s}}\mid s\in S\}$.
@@
$\Delta(k;\bm{s})=\Delta_{w\varpi_{s},w^{\prime}\varpi_{s}}.$
@@
and do the same procedure for $w^{\prime}$. Moreover, if $w=e$ (resp. $w^{\prime}=e$), there is no condition on $(s_{i_{1}},\dots,s_{i_{N/2}})$ (resp. $(s_{j_{1}},\dots,s_{j_{N/2}})$). Set
@@
Note that neither $w_{0}\varpi_{s}$ nor $w\varpi_{s}$ with $w\varpi_{s}<\varpi_{s}$ is dominant.
@@ 3.1. Main result
We do not state that the statements of Theorems 3.1 and 3.2 do not hold in the case when $G$ is of type $F_{4}$. Indeed, $\mathbb{C}[G]=\mathscr{U}$ is proved for an arbitrary $G$ in a very recent preprint [QY25, Theorem B.1] by a different method.
@@
It suffices to show that $C^{V}(f,v)$ can be expressed as a polynomial in generalized minors for weight vectors $f\in(V^{\ast})_{\mu}$ and $v\in V_{\nu}$. If $\mu,\nu\in\operatorname{wt}(V)\setminus\{0\}$, then $C^{V}(f,v)$ is a generalized minor up to a constant. Hence we deal with the case when $\mu$ or $\nu$ equals 0.
@@
for some $v^{\prime}_{\eta}\in V_{\eta}$ and $v^{\prime\prime}_{-\eta}\in V_{-\eta}$. Hence, if we write $m^{\ast}(f)$ as $\sum_{i\in I}f^{\prime}_{i}\otimes f^{\prime\prime}_{i}$ by some $f^{\prime}_{i},f^{\prime\prime}_{i}\in V^{\ast}$ and a finite index set $I$, then
@@
$=C^{V\otimes V}\left(\sum\nolimits_{i\in I}f^{\prime}_{i}\otimes f^{\prime\prime}_{i},\sum\nolimits_{\eta\in P;\eta\neq 0}v^{\prime}_{\eta}\otimes v^{\prime\prime}_{-\eta}\right)$
$=\sum_{i\in I}\left(\sum_{\eta\in P;\eta\neq 0}C^{V}(f^{\prime}_{i},v^{\prime}_{\eta})C^{V}(f^{\prime\prime}_{i},v^{\prime\prime}_{-\eta})\right).$
@@
When $\mathfrak{g}$ is of type $E_{8}$ or $G_{2}$, there exist a quasi-minuscule $\mathfrak{g}$-module $V$ and a surjective homomorphism $m\colon V\otimes V\to V$ of $\mathfrak{g}$-modules satisfying $m(V_{0}\otimes V_{0})=\{0\}$.
@@
Let $\mathfrak{g}$ be a simple Lie algebra of type $E_{8}$ or $G_{2}$. Consider the quasi-minuscule $\mathfrak{g}$-module $V$ stated in Theorem 3.9. Then the corresponding representation of $G$ is a faithful representation $\rho\colon G\to GL(V)$ which gives a closed immersion $G\to SL(V)$, since the center of $G$ is trivial in this case (see [Bor91, Lemma 22.1], [GW09, Remark 11.2.16], [Bri10, Corollary 1.13]). Therefore, the coordinate ring $\mathbb{C}[G]$ is generated by the matrix coefficients of $V$ (see, for example, [IOS23, Proposition 2.1]), and hence, by the generalized minors thanks to Theorem 3.8. ∎
@@
$m\colon V(\varpi_{2})\otimes V(\varpi_{2})\hookrightarrow\mathfrak{C}\otimes\mathfrak{C}\xrightarrow{m_{\mathfrak{C}}}\mathfrak{C}\simeq V(\varpi_{2})\oplus V(0)\xrightarrow{\text{projection}}V(\varpi_{2}).$
When $\mathfrak{g}$ is of type $F_{4}$, there does not exist a homomorphism $m$ of $\mathfrak{g}$-modules satisfying the assumption in Theorem 3.8. Indeed, a quasi-minuscule $\mathfrak{g}$-module is isomorphic to the 26-dimensional irreducible module $V(\varpi_{4})$ ($\dim V(\varpi_{4})_{0}=2$). We have $V(\varpi_{4})\otimes V(\varpi_{4})\simeq V(2\varpi_{4})\oplus V(\varpi_{1})\oplus V(\varpi_{3})\oplus V(\varpi_{4})$. Therefore, a surjective homomorphism of $\mathfrak{g}$-modules $m\colon V(\varpi_{4})\otimes V(\varpi_{4})\to V(\varpi_{4})$ is uniquely determined up to constant multiplication, and it is the only candidate for the homomorphism $m$ satisfying the assumption in Theorem 3.8. However, we have $m(V(\varpi_{4})_{0}\otimes V(\varpi_{4})_{0})\neq\{0\}$. The Lie algebra $\mathfrak{g}$ of type $F_{4}$ is realized as the Lie algebra of derivations of the $27$-dimensional exceptional Jordan algebra $\mathfrak{J}$ over $\mathbb{C}$. See, for example, [Yok25, Section 2] for its explicit presentation. Hence there is an action of $\mathfrak{g}$ on $\mathfrak{J}$, and $\mathfrak{J}$ is isomorphic to $V(\varpi_{4})\oplus V(0)$ as a $\mathfrak{g}$-module. Here the trivial $\mathfrak{g}$-module $V(0)$ corresponds to the space spanned by the unit $\mathbf{1}$ of $\mathfrak{J}$. Then the multiplication $m_{\mathfrak{J}}$ of $\mathfrak{J}$ induces a surjective homomorphism of $\mathfrak{g}$-modules $m\colon V(\varpi_{4})\otimes V(\varpi_{4})\to V(\varpi_{4})$ as follows:
$m\colon V(\varpi_{4})\otimes V(\varpi_{4})\hookrightarrow\mathfrak{J}\otimes\mathfrak{J}\xrightarrow{m_{\mathfrak{J}}}\mathfrak{J}\simeq V(\varpi_{4})\oplus V(0)\xrightarrow{\text{projection}}V(\varpi_{4}).$
However, we can directly check that $m(V(\varpi_{4})_{0}\otimes V(\varpi_{4})_{0})\neq\{0\}$ since the weight space $V(\varpi_{4})_{0}$ corresponds to the space spanned by the trace-free diagonal matrices in $\mathfrak{J}$ under the presentation in [Yok25, Section 2]. See [Yok25] for further details.